The critical angle at an air-liquid interface is 50°. Calculate the refractive index of the liquid

1
Expert's answer
2022-02-08T10:02:10-0500

Given:

"\theta_c=50^\circ"


The critical angle is given by

"\sin\theta_c=\frac{1}{n}"

Hence, he refractive index of the liquid

"n=\frac{1}{\sin\theta_c}=\frac{1}{\sin50^\circ}=1.3"